OpenStudy (anonymous):

Which of the following is an odd function? f(x) = 3x2 + x f(x) = 4x3 + 7 f(x) = 5x2 + 9 f(x) = 6x3 + 2x

OpenStudy (anonymous):

\[a \ function \ f(x) \ is \ called\ an\ odd\ function\ if\ f(-x)=-f(x)\\ clearly\ last\ one\ f(x)=6x^3+2x\ satisfies\ this\ definition\\since\ f(-x)=6(-x)^3+2(-x)\\ f(-x)=-6x^3-2x\\f(-x)= -[6x^3+2x]\\f(-x)=-f(x)\\ \therefore\ f(x)= 6x^3+2x\ is\ an\ odd\ function'\]

OpenStudy (anonymous):

ohhh okay thank you!!!

OpenStudy (anonymous):

cheers!